# Average Certified Radius is a Poor Metric for Randomized Smoothing

### v2

@@ 8 Conclusion
This work rigorously demonstrates that Average Certified Radius (ACR) is a poor metric for Randomized Smoothing (RS). Theoretically, we prove that ACR of a trivial classifier can be arbitrarily large, and that an improvement on easy inputs contributes much more to ACR than on hard inputs. Empirically, we show that all state-of-the-art (SOTA) strategies reduce the accuracy on hard inputs and only focus on easy inputs to increase ACR. Based on these novel insights, we develop strategies to amplify Gaussian training by reweighting the training data to focus only on easy inputs. Specifically, we discard hard inputs during training, weight the dataset with their contribution to ACR, and apply extreme optimization for easy inputs via adversarial noise selection. With these intuitive modifications to the simple Gaussian training, we replicate the effect of SOTA training algorithms and achieve the SOTA ACR. Overall, our results suggest the need for evaluating RS training with better metrics. To this end, we suggest using the best certified accuracy at various radii and reporting the empirical distribution of $p_{A}$ for a large $N$ in replacement of ACR. We hope this work will inspire future research in this direction and motivate the community to permanently abandon ACR in the evaluation of randomized smoothing.
@@ A.2 Proof of Theorem 2
Let $\delta=[q_{1},q_{2},\dots,q_{d}]\in{\mathbb{R}}^{d}$ be sampled from ${\mathcal{N}}(0,\sigma^{2}I_{d})$. Then we have
$\mathbb{P}(\delta)$ $=\frac{1}{(2\pi)^{d/2}\sigma^{d}}\exp\left(-\frac{1}{2\sigma^{2}}\sum_{i=1}^{d}q_{i}^{2}\right)$
$=\frac{1}{(2\pi)^{d/2}\sigma^{d}}\exp\left(-\frac{1}{2\sigma^{2}}\|\delta\|_{2}^{2}\right).$
@@ Appendix B On the Empirical Distribution of $p_{A}$
In this section, we include a case study to further demonstrate the idea we presented in Section 7 about using the empirical distribution of $p_{A}$ to evaluate the RS model. We first show how to calculate the certified accuracy for every interested radius and certification budget efficiently, given the ECDF of $p_{A}$. Then, we show how to convert an existing certified accuracy-radius curve to the ECDF of $p_{A}$, thus the literature results can be easily reused. These features facilitate practitioners to apply RS models in practice, as the certification budget usually varies in different scenarios.
@@ B.1 From ECDF to Certified Accuracy under Different Certification Settings
We consider four RS models trained on CIFAR-10. The ECDF of $p_{A}$ under $N=10^{5}$ is shown in Figure 7. Assume now instead of certifying with $N=10^{5}$ and $\alpha=0.001$, we want to certify with $N=100$ and $\alpha=0.01$, which is less expensive and more practical to run on edge devices. We are interested in the certified accuracy-radius curve under this new certification setting.
@@ C.1 Training Details
We follow the standard training protocols in previous works. Specifically, we use ResNet-110 (He et al., 2016) for CIFAR-10. We investigate three noise levels, $\sigma=0.25,0.5,1.0$. We train the model for 150 epochs with an initial learning rate of 0.1, which is decreased by a factor of 10 in every 50 epochs. Stochastic gradient descent (SGD) optimizer with a momentum of 0.9 is used, and the batch size is set to 256. We adjust the following hyperparameters to find the best performance for each setting:
@@
All other parameters are fixed to their default values. Specifically, we use 100 noise samples ($N=100$) to calculate $p_{A}$ when discarding hard samples. For updating dataset weight, we use 16 noise samples ($N=16$) to calculate $p_{A}$, with a minimum value set to 0.75 ($p_{\rm min}$=0.75). The number of noise samples for each input is set to $m=4$. The hyperparameters we use on CIFAR-10 are shown in Table 4.
@@
$\sigma$ 0.25 0.5 1.0 $E_{t}$ 60 70 60 $p_{t}$ 0.5 0.4 0.4 $T$ 3 6 4 $\epsilon$ 0.25 0.25 0.5
Table 4: Hyperparameters we use on CIFAR-10.
C.2 Adversarial Attack Algorithm
@@
where $L$ is the loss function, $f$ is the model, $x$ is the input, $y$ is the label, $\delta$ is the noise, $\delta^{*}$ is the updated noise, sign is the sign function, and $\epsilon$ is the step size. Without specific instructions, our experiments are always conducted with $l_{2}$ PGD. We show the results with $l_{\infty}$ PGD on CIFAR-10 in Section E.3.
@@
Table 5: Mean and Standard deviation of ACR and certified radius on CIFAR-10 using our method. The results are based on three independent runs with different random seeds.
@@ Appendix D Additional Ablation Studies
In this section, we provide additional ablation studies on the effect of different hyperparameters. All results are based on $\sigma=0.5$ and CIFAR-10. Unless otherwise specified, all hyperparameters are the same as those in Table 4.
@@ D.2 Effect of Adaptive Attack
(a) Effect of $\epsilon$
(b) Effect of $T$
@@
Appendix E Additional Results
E.1 Result Significance
Table 5 shows the mean and standard deviation of the ACR and certified accuracy-radius curve on CIFAR-10 using our best hyperparameters reported in Table 4. The results are based on three independent runs with different random seeds. The standard deviation is relatively small, indicating that the results are stable and reproducible
E.2 Discarded Data Ratio
The remaining data ratios after discarding hard samples under different $\sigma$ on CIFAR-10 are $91\%$, $80\%$ and $58\%$, when $\sigma=0.25$, $0.5$ and $1.0$ respectively. The results indicate that even though a significant portion of the training set is discarded during training, we still achieve a better ACR than the current SOTA model according to Table 2. This indicates that discarding a large amount of data during training leads to a better ACR, proving once again that ACR may not accurately represent the overall robustness of the model.
E.3 Replacing $l_{2}$ with $l_{\infty}$ PGD Attack
In this section, we show the results under $l_{\infty}$ PGD attack in Table 10 and Table 11. For $l_{\infty}$ attack, we use much smaller step sizes $\epsilon$ to have the similar attack strength to $l_{2}$ attack. Specifically, we investigate $\sqrt{d}\cdot\epsilon=0.125,0.25,1.5,1.0$ for $l_{\infty}$ attack, where $d$ is the dimension of the input. After the thorough experiments, we find that using $\sqrt{d}\cdot\epsilon=0.25$ and the same with $l_{2}$ attack for other hyperparameters reaches the best ACR.
The $l_{\infty}$ attack has the similar effect on results to the $l_{2}$ attack. Increasing the step size or the number of steps of the attack will result in better accuracy at large radii but lower accuracy at small radii.

### v4

@@ 8 Conclusion
This work rigorously demonstrates that Average Certified Radius (ACR) is a poor metric for Randomized Smoothing (RS). Theoretically, we prove that ACR of a trivial classifier can be arbitrarily large, and that an improvement on easy inputs contributes much more to ACR than the same amount of improvement on hard inputs. In addition, the comparison results based on ACR varies when the certification budget changes, thus unsuitable for comparing algorithms. Empirically, we show that all state-of-the-art (SOTA) strategies reduce the accuracy on hard inputs and only focus on easy inputs to increase ACR. Based on these novel insights, we develop strategies to amplify Gaussian training by focusing only on easy data points during training. Specifically, we discard hard inputs during training, reweigh the dataset with their contribution to ACR, and apply extreme optimization for easy inputs via adversarial noise selection. With these intuitive modifications to the simple Gaussian training, we replicate the effect of SOTA training algorithms and achieve the SOTA ACR. Overall, our results suggest the need for evaluating RS training with better metrics. To this end, we suggest reporting the empirical distribution of $p_{A}$ for a large $N$ in replacement of ACR. We hope this work will inspire future research in this direction and motivate the community to permanently abandon ACR in the evaluation of randomized smoothing.
Reproducibility Statement
Experimental details, including hyperparameters, datasets, and training procedures, are provided in App. C. Result significance is confirmed in Section E.1. Code and models are available at https://github.com/eth-sri/acr-weakness.
Acknowledgements
This work has been done as part of the EU grant ELSA (European Lighthouse on Secure and Safe AI, grant agreement no. 101070617) and the SERI grant SAFEAI (Certified Safe, Fair and Robust Artificial Intelligence, contract no. MB22.00088). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or European Commission. Neither the European Union nor the European Commission can be held responsible for them.
The work has received funding from the Swiss State Secretariat for Education, Research and Innovation (SERI).
@@ A.2 Proof of Theorem 2
Let $\bm{\delta}=[q_{1},q_{2},\dots,q_{d}]\in{\mathbb{R}}^{d}$ be sampled from ${\mathcal{N}}(0,\sigma^{2}I_{d})$. Then we have
$\mathbb{P}(\bm{\delta})$ $=\frac{1}{(2\pi)^{d/2}\sigma^{d}}\exp\left(-\frac{1}{2\sigma^{2}}\sum_{i=1}^{d}q_{i}^{2}\right)$
$=\frac{1}{(2\pi)^{d/2}\sigma^{d}}\exp\left(-\frac{1}{2\sigma^{2}}\|\bm{\delta}\|_{2}^{2}\right).$
@@ Appendix B On the Empirical Distribution of $p_{A}$
In this section, we include a case study to further demonstrate the idea we presented in Section 7 about using the empirical distribution of $p_{A}$ to evaluate the RS model. We first show how to calculate the certified accuracy for every interested radius and certification budget efficiently, given the ECDF of $p_{A}$. Then, we show how to convert an existing certified accuracy-radius curve to the ECDF of $p_{A}$, thus the literature results can be easily reused. These features facilitate practitioners to apply RS models in practice, as the certification budget usually varies in different scenarios. Finally, we re-evaluate previous RS training strategies with the ECDF of $p_{A}$.
@@ B.1 From ECDF to Certified Accuracy under Different Certification Settings
We consider four RS models (Gaussian, Consistency, CAT-RS and ours) trained on CIFAR-10. The ECDF of $p_{A}$ under $N=10^{5}$ is shown in Figure 7. Assume now instead of certifying with $N=10^{5}$ and $\alpha=0.001$, we want to certify with $N=100$ and $\alpha=0.01$, which is less expensive and more practical to run on edge devices. We are interested in the certified accuracy-radius curve under this new certification setting.
@@ B.2 From Certified Radius to ECDF
B.3 Evaluating Current RS with ECDF
For future reference, we evaluate all existing methods, including Gaussian, Consistency, SmoothAdv, SmoothMix, MACER, CAT-RS and ours, using the ECDF of $p_{A}$, together with our proposed algorithm. We use the same dataset and model as in Section 6, i.e. ResNet-110 on CIFAR-10 with $\sigma\in\{0.25,0.5,1.0\}$ and $N=10^{5}$. Whenever an official public model is available, we use it to compute the ECDF of $p_{A}$. Otherwise, we re-train the model with the same hyperparameters as in the original paper.
The ECDF of $p_{A}$ is shown in Figure 9. The related data point is available in our GitHub repository. We note that the ECDF is independent of $\alpha$, thus it can be used to evaluate the certified accuracy under any certification setting. Given the ECDF metric, model ${\mathcal{A}}$ is better than model ${\mathcal{B}}$ if and only if ${\mathcal{A}}$ has higher ECDF than ${\mathcal{B}}$ for all $p_{A}\geq 0.5$. Therefore, it is possible that neither of two models is better than the other.
For various $\sigma$, the Gaussian model is always the best in the low $p_{A}$ region, i.e. from $p_{A}=0.5$ to around 0.65. Although the our model consistently has the highest ACR under all $\sigma$, it only outperforms other models in the high $p_{A}$ region. For example, with $\sigma=0.5$, our model is the worst when $p_{A}$ is below 0.75. We list some cases where two models have strict orders: with $\sigma=0.25$, the CAT-RS model is better than Consistency and SmoothAdv, and our model is better than CAT-RS; with $\sigma=0.5$, the CAT-RS model is better than the SmoothMix model; with $\sigma=1.0$, the CAT-RS model is better than SmoothMix and MACER, and our model is better than MACER.
(a) $\sigma=0.25$
(b) $\sigma=0.5$
(c) $\sigma=1.0$
Figure 9: The empirical distribution of $p_{A}$ on CIFAR-10 with different $\sigma$ and $N=10^{5}$. This figure is an extension of Figure 7 on more $\sigma$ and algorithms.
@@ C.1 Training Details
We follow the standard training protocols in previous works. Specifically, we use ResNet-110 (He et al., 2016) on CIFAR-10 and ResNet-50 on ImageNet, respectively. For CIFAR-10, we train ResNet-110 for 150 epochs with an initial learning rate of 0.1, which is decreased by a factor of 10 in every 50 epochs. The inital learning rate of training ResNet-50 on ImageNet is 0.1 and it is decreased by a factor of 10 in every 30 epochs within 90 epochs. We investigate three noise levels, $\sigma=0.25,0.5,1.0$. We use SGD with momentum 0.9. The following hyperparameter is tuned for each setting:
[…]
@@
All other hyperparameter is fixed to the default values. Specifically, for CIFAR-10 we use 100 noise samples to calculate $p_{A}$ when discarding hard samples. When updating dataset weight, we use 16 noise samples to calculate $p_{A}$, with $p_{\rm min}=0.75$. The hyperparameter tuned on CIFAR-10 is shown in Table 6. For ImageNet, we skip the ‘Data Reweighting with Certified Radius’ step for faster training, and use a Gaussian-pretrained ResNet-50 to evaluate the $p_{A}$ with 50 noise samples. The hyperparameters tuned on ImageNet is shown in Table 7.
$\sigma$ 0.25 0.5 1.0 $m$ 4 4 4 $E_{t}$ 60 70 60 $p_{t}$ 0.5 0.4 0.4 $T$ 3 6 4 $\epsilon$ 0.25 0.25 0.5
Table 6: Hyperparameters we use on CIFAR-10.
$\sigma$ 0.25 0.5 1.0 $m$ 1 1 2 $E_{t}$ 0 0 0 $p_{t}$ 0.2 0.2 0.1 $T$ 2 2 1 $\epsilon$ 0.5 1.0 2.0
Table 7: Hyperparameters we used on ImageNet.
In Algorithm 2, we present a pseudocode for the overall training procedure described in Section 5.
Algorithm 2 Overall Training Procedure
  for epoch $=1$ to $N_{\rm epoch}$ do
[…]
  end if
[…]
  end for
  end for
[…]
C.2 Certification Algorithm
@@
C.3 Adversarial Attack Algorithm
@@
where $L$ is the loss function, $f$ is the model, $x$ is the input, $y$ is the label, $\delta$ is the noise, $\delta^{*}$ is the updated noise, sign is the sign function, and $\epsilon$ is the step size. Unless stated otherwise, our experiments are always conducted with the $l_{2}$ PGD attack. We show the results with the $l_{\infty}$ PGD attack on CIFAR-10 in Section E.4.
(a) Effect of $\epsilon$
(b) Effect of $T$
(a) $\sigma=0.25$
(b) $\sigma=0.5$
(c) $\sigma=1.0$
Figure 11: The empirical cumulative distribution (ECDF) of $p_{A}$ on ImageNet for models trained and certified with various $\sigma$ with different training algorithms.
@@
Table 8: Mean and Standard deviation of ACR and certified accuracy on CIFAR-10 using our method. The results are based on three independent runs with different random seeds.
@@ Appendix D Additional Ablation Studies
In this section, we provide additional ablation studies on the effect of different hyperparameters. All results are based on $\sigma=0.5$ and CIFAR-10. Unless stated otherwise, all hyperparameters are the same as those in Table 6.
@@ D.1 Effect of Discarding Hard Data
D.2 Effect of Adaptive Attack
Appendix E Additional Results
E.1 Result Significance
Table 8 shows the mean and standard deviation of the ACR and certified accuracy-radius curve on CIFAR-10 using our best hyperparameters reported in Table 6. The results are based on three independent runs with different random seeds. The standard deviation is relatively small, confirming the stability of our results.
E.2 Empirical Distribution of $p_{A}$ on ImageNet
We extend Figure 3 to ImageNet, shown as Figure 11. We observe that for SOTA methods, there are more samples with $p_{A}$ below 0.5 compared to Gaussian training, which is consistent to the results on CIFAR-10 in Section 4.3.
E.3 Discarded Data Ratio
In this section, we are interested in how many data samples are discarded during training. The remaining data ratios under different $\sigma$ on CIFAR-10 are $91\%$, $80\%$ and $58\%$, for $\sigma=0.25$, $0.5$ and $1.0$, respectively. The results indicate that even though a significant portion of the training set is discarded during training, we still achieve a better ACR than the current SOTA model according to Table 3. Further, this confirms that larger $\sigma$ leads to more discarded data samples, which is consistent with the intuition.
E.4 Replacing $l_{2}$ with $l_{\infty}$ PGD Attack
@@
In this section, we replace the $L_{2}$ attack with $l_{\infty}$ PGD attack. The results are shown in Table 13 and Table 14. For $l_{\infty}$ attack, we use much smaller step sizes $\epsilon$ to have similar attack strengths to $l_{2}$ attack. Specifically, we investigate $\sqrt{d}\cdot\epsilon=0.125,0.25,1.5,1.0$ for $l_{\infty}$ attack, where $d$ is the dimension of the input. After the thorough experiments, we find that using $\sqrt{d}\cdot\epsilon=0.25$ with other hyperparameters fixed to the values in Table 6 has the best ACR. The $l_{\infty}$ attack has a similar effect to the $l_{2}$ attack. Increasing the step size and the number of steps of the attack result in better accuracy at large radii but lower accuracy at small radii.